OpenStudy (anonymous):

Some basic algebraic concepts: a(b+c) = ? Thus, 10(x+2) = ?

OpenStudy (anonymous):

a(b+c) = ab + ac

HanAkoSolo (jamierox4ev3r):

wouldn't you just distribute it?

HanAkoSolo (jamierox4ev3r):

10x+20

OpenStudy (anonymous):

10(x+2) = 10*x + 10*2 = ____

OpenStudy (anonymous):

@Jamierox4ev3r wouldn't it be better if we tried to not just give out the answer? :) thank you.